# Networked Communication for Decentralised Cooperative Agents in Mean-Field Control

### v1

@@ 4 Theoretical results
In coordination and anti-coordination games where Ass. 1, 2, 3, 4, 5 and 6 apply, we have $\mathbb{E}[W(\pi^{\mathrm{net}}_{k+1},I(\pi^{\mathrm{net}}_{k+1}))]>\mathbb{E}[W(\pi^{\mathrm{cent}}_{k+1},I(\pi^{\mathrm{cent}}_{k+1}))]$ (i.e. in expectation networked agents will increase their returns faster than centralised ones).
Proof sketch. A policy that successfully spreads through a networked population based on a metric relating to its performance is expected to have better base policy quality $b(\pi)$ than that of a policy arbitrarily forced on a population by a central learner. Full proof in Appx. C.1.
@@
We now give results showing why learning can be faster in the networked case than the independent case. However, since we cannot expect independent agents to share a single policy $\pi_{k+1}$ after the update in each iteration, it is not possible to extract a solution to the MFC problem from the independent policies (a further weakness of the independent case). We therefore give these results in terms of the population-average return (Def. 3) instead of the social welfare (Def. 5) as before.
Assumption 7.
@@
Assume that the estimated global mean field and average reward in the networked case are the same as the independent case, i.e. $\forall i,j\;\tilde{\hat{\mu}}^{(i,net)}_{t}$ = $\tilde{\hat{\mu}}^{(j,ind)}_{t}$ and $\tilde{\hat{r}}^{(i,net)}_{t}=r^{i}_{t}$. [^5] In other words, we assume for simplicity that the only difference between the networked and independent cases is the networked policy communication scheme. In practice, the networked estimates will be better due to communication, giving an additional performance increase over the independent case.
@@
We refer to the joint policy in the networked case after communication round $c$ as $\boldsymbol{\pi}^{\mathrm{net}}_{k+1,c}=\left(\pi^{(1,\mathrm{net})}_{k+1,c},\dots,\pi^{(N,\mathrm{net})}_{k+1,c}\right)$, and the joint policy in the networked case as $\boldsymbol{\pi}^{\mathrm{ind}}_{k+1}=\left(\pi^{(1,\mathrm{ind})}_{k+1},\dots,\pi^{(N,\mathrm{ind})}_{k+1}\right)$.
Theorem 2.
In a coordination game, given Ass. 2, 3, 4 and 7, even a single round of communication in the networked case improves on the independent case, i.e. for $c=0$, $\mathbb{E}\left[V^{pop}(\boldsymbol{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{0})\right]>\mathbb{E}\left[V^{pop}(\boldsymbol{\pi}^{\mathrm{ind}}_{k+1},\mu_{0})\right]$.
Proof sketch. If agents replace their policies with neighbours’ policies that are expected to have better base quality $b(\pi)$ than their own, then both the average base quality and the average alignment should improve, leading to a better return in coordination games. Full proof in Appx. C.2.
To prove the benefit of the networked case over the independent case in anti-coordination games, we use an additional assumption.
Assumption 8.
@@
Assume that an increase in the base return function outweighs a decrease in the population’s policy diversity, namely $h(b+\Delta b,f_{d}-\Delta f_{d})>h(b,f_{d}),\;\forall\Delta b>0,\Delta f_{d}>0.$
Theorem 3.
In an anti-coordination game, given Ass. 2, 3, 4, 7 and 8, even a single round of communication in the networked case improves on the independent case, i.e. for $c=0$, $\mathbb{E}\left[V^{pop}(\boldsymbol{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{0})\right]>\mathbb{E}\left[V^{pop}(\boldsymbol{\pi}^{\mathrm{ind}}_{k+1},\mu_{0})\right]$.
Proof sketch. If agents replace their policies with neighbours’ policies that are expected to have better base quality $b(\pi)$ than their own, then the average return will increase, assuming the increase in base quality outweighs the decrease in diversity. Full proof in Appx. C.3.
@@ 5.1 Experimental setup
For the types of game used in our experiments we follow the gold standard in similar works on MFGs, namely grid-world environments where agents can move in the four cardinal directions or remain in place (Laurière, 2021; Laurière et al., 2022b; Zaman et al., 2023; Algumaei et al., 2023; Cui et al., 2023a; Benjamin & Abate, 2023; 2024; Wu et al., 2024). We present results from six tasks similar to those found in prior works, defined by the agents’ reward/transition functions. Two are coordination games and four are anti-coordination games, where in each case the reward function reflects a coordination/anti-coordination ($f_{c}/f_{d}$) element alongside other elements that may be crucial for receiving reward, reflected in the policies’ base quality $b(\pi)$ (Sec. 4). See Appx. E.1 for full technical descriptions of the tasks. The two coordination games are:
  • Cluster. Agents are rewarded for gathering but must agree where to do so among themselves.
  • Target selection. Agents are rewarded for visiting any of a given number of targets, but the reward is proportional to the number of other agents co-located at the target. Agents must coordinate on which single target they will all meet at to maximise their individual rewards.
@@
  • Disperse. Agents are rewarded for being located in more sparsely populated areas, but only if they are stationary (to discourage trivial random policies).
  • Target coverage. The population is rewarded for spreading across a certain number of targets, as long as agents are stationary at the target.
  • Beach bar. Agents are rewarded for being stationary in sparsely populated locations as close as possible to a target.
  • Shape formation. The population is rewarded for spreading around a ring shape.
In these spatial environments, the communication network $\mathcal{G}^{comm}_{t}$ is determined by the physical distance from agent $i$; we show plots for various broadcast radii, expressed as fractions of the maximum possible distance (the grid’s diagonal length). We evaluate our experiments according to a finite-step approximation of the discounted population-average return (Def. 3) over $M$ steps within each outer $k$ loop, i.e. $\hat{V}^{pop}(\boldsymbol{\pi}_{k},\mu_{0};M)$. We discuss hyperparameters in Appx. E.2.

### v2

@@ 5.2 Analysis
Let us set $\tau^{comm}_{k}\in\mathbb{R}_{>0}$. In coordination and anti-coordination games where Assumptions 5.4-5.8 apply, we have $\mathbb{E}[W(\pi^{\mathrm{net}}_{k+1},I(\pi^{\mathrm{net}}_{k+1}))]\geq\mathbb{E}[W(\pi^{\mathrm{cent}}_{k+1},I(\pi^{\mathrm{cent}}_{k+1}))]$.
Remark 5.10.
Assumption 5.5 presumes that all policies are pairwise distinct after the updates, but does not restrict their returns in the same way. If we additionally make the very weak assumption that at least one of these distinct policies in each $k$ iteration has a base return that is distinct from the others (which is likely to hold in all but the most trivial environments), the inequality in the theorem above will be strict, i.e. networked learning will always be faster in expectation.
Proof.
Recall that before the communication rounds in Line 8 (Alg. 3), the randomly updated policies $\pi^{i}_{k+1}$$_{i=1}^{N}$ have associated estimated returns $\sigma^{i}_{k+1}$$_{i=1}^{N}$. Denote the mean and maximum of this set $\sigma^{\mathrm{mean}}_{k+1}$ and $\sigma^{\mathrm{max}}_{k+1}$ respectively. Since $\pi^{\mathrm{cent}}_{k+1}$ is chosen arbitrarily from $\pi^{i}_{k+1}$$_{i=1}^{N}$, it will obey $\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]=\sigma^{\mathrm{mean}}_{k+1}$ $\forall k$, though there will be high variance. Conversely, for the networked case the softmax adoption scheme (Line 11, Alg. 3), which for $\tau^{comm}_{k}\in\mathbb{R}_{>0}$ gives non-uniform adoption probabilities for distinct $\sigma$ values, means by definition that some communicated policies are more likely to be adopted than others if they have distinct finitely estimated returns (those with higher $\sigma^{i}_{k+1}$ are more likely to be adopted at each communication round). Thus the consensus $\pi^{\mathrm{net}}_{k+1}$ that gets adopted by the whole networked population will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\sigma^{\mathrm{mean}}_{k+1}$ if at least one policy receives a distinct return from the others, or $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]\geq\sigma^{\mathrm{mean}}_{k+1}$ in the rare circumstance that all policies receive the same return. If $\tau^{comm}_{k+1}\rightarrow 0$, it will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]=\sigma^{\mathrm{max}}_{k+1}$ $\forall k$. As such:
$\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]\geq\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}].$ (1)
In Eq. 1 and the remaining equations of the proof, bear in mind that the equality will be strict if at least one policy receives a distinct return from the others.
Refer to the agent whose update originally gave rise to $\pi^{\mathrm{net}}_{k+1}$ and $\sigma^{\mathrm{net}}_{k+1}$ as agent $(i,\mathrm{net})$; we equivalently also have the arbitrary agent $(j,\mathrm{cent})$. Prior to consensus being attained in each case, the joint policy can be written as $\bm{\pi}^{(i,\mathrm{net};j,\mathrm{cent})}$ $(\pi^{1},\dots,\pi^{i-1},\pi^{(i,\mathrm{net})},\pi^{i+1},\dots,\pi^{j-1},\pi^{(j,\mathrm{cent})},\pi^{j+1},\dots,$ $\pi^{N})$.
Given Eq. 1, and by Assumption 5.6 on the quality of finite-step estimations, we know that directly after the policy update in Line 16 (Alg. 2), prior to the consensus being reached, we have:
$\mathbb{E}\left[V^{(i,\mathrm{net})}(\bm{\pi}^{(i,\mathrm{net};j,\mathrm{cent})}_{k+1},\mu_{t})\right]\geq\mathbb{E}\left[V^{(j,\mathrm{cent})}(\bm{\pi}^{(i,\mathrm{net};j,\mathrm{cent})}_{k+1},\mu_{t})\right].$ (2)
By Assumption 5.5 we know that straight after the random policy updates there is no alignment among policies, i.e. in a coordination game we have $f_{c}^{(i,\mathrm{net})}$ = $f_{c}^{(j,\mathrm{cent})}=\min f_{c}$, and in an anti-coordination game we have $f_{d}^{(i,\mathrm{net})}$ = $f_{d}^{(j,\mathrm{cent})}=\max f_{d}$. Therefore if Eq. 2 pertains, by Def. 5.2 it must be because:
$\mathbb{E}[b(\pi^{(i,\mathrm{net})})]\geq\mathbb{E}[b(\pi^{(j,\mathrm{cent})})],$ (3)
By Assumption 5.7 on policy consensus, we know that in the networked and central-agent cases the joint policies respectively become $\bm{\pi}^{\mathrm{net}}:=(\pi^{\mathrm{net}},\pi^{\mathrm{net}},\pi^{\mathrm{net}},\dots)$ and $\bm{\pi}^{\mathrm{cent}}:=(\pi^{\mathrm{cent}},\pi^{\mathrm{cent}},\pi^{\mathrm{cent}},\dots)$. We therefore end up with maximum alignment in both cases, such that $f_{c}^{\mathrm{net}}$ = $f_{c}^{\mathrm{cent}}=\max f_{c}$ in a coordination game, and $f_{d}^{\mathrm{net}}$ = $f_{d}^{\mathrm{cent}}=\min f_{d}$ in an anti-coordination game. Due to this, along with Eqs. 2 and 3, we have
$\mathbb{E}\left[V^{i}(\bm{\pi}^{\mathrm{net}}_{k+1},\mu_{t})\right]\geq\mathbb{E}\left[V^{j}(\bm{\pi}^{\mathrm{cent}}_{k+1},\mu_{t})\right].$ (4)
In turn we have:
$\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1},\mu_{t})\right]\geq\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{cent}}_{k+1},\mu_{t})\right],$ (5)
which by Assumption 5.8 gives
$\mathbb{E}[W(\pi^{\mathrm{net}}_{k+1},I(\pi^{\mathrm{net}}_{k+1}))]\geq\mathbb{E}[W(\pi^{\mathrm{cent}}_{k+1},I(\pi^{\mathrm{cent}}_{k+1}))],$
namely the result. ∎
@@
We now give results showing that learning is at least as fast in the networked case than in the independent case - empirically we find networked learning always to be strictly faster. We give separate theorems for coordination and anti-coordination games. Since we cannot necessarily expect the independent agents to share a single policy $\pi_{k+1}$ after the update in each iteration of learning, we give these results in terms of the population-average return (Def. 3.3) instead of the single-policy social welfare (Def. 3.5) as before.
@@
Again, we assume for simplicity of the theory that it is only the policy communication scheme that creates a difference in learning between the networked and independent cases, i.e. we assume that networked agents receive the same estimates of the mean field and average reward as independent agents. As mentioned above, our ablation studies suggest this is the dominant factor in our experimental settings anyway, with the estimated mean field not required at all in the broad class of stationary problems. Nevertheless, in practice the networked estimates of the (mean field and) average reward will be better than the independent ones, giving an additional performance increase over the independent case. Thus loosening this assumption is likely to actually enhance the effects identified in the theorems.
Assumption 5.11.
Assume that the estimated global mean field and average reward in the networked case are the same as the independent case, i.e. $\forall i,j\;\tilde{\hat{\mu}}^{(i,net)}_{t}$ = $\tilde{\hat{\mu}}^{(j,ind)}_{t}$ and $\tilde{\hat{r}}^{(i,net)}_{t}=r^{i}_{t}$.
@@
We refer to the joint policy in the networked case after communication round $c$ as $\bm{\pi}^{\mathrm{net}}_{k+1,c}=\left(\pi^{(1,\mathrm{net})}_{k+1,c},\dots,\pi^{(N,\mathrm{net})}_{k+1,c}\right)$, and the joint policy in the independent case as $\bm{\pi}^{\mathrm{ind}}_{k+1}=\left(\pi^{(1,\mathrm{ind})}_{k+1},\dots,\pi^{(N,\mathrm{ind})}_{k+1}\right)$.
We can now give our second theorem, namely that in expectation networked populations will increase their returns at least as fast as independent ones in coordination games with only a single round of communication in each iteration.
Theorem 5.12.
Let us again set $\tau^{comm}_{k}\in\mathbb{R}_{>0}$. In a coordination game, given Assumptions 5.5, 5.6 and 5.11, for $c=0$, $\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})\right]\geq\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{ind}}_{k+1},\mu_{t})\right]$.
Remark 5.13.
Assumption 5.5 presumes that all policies are pairwise distinct after the updates, but does not restrict their returns in the same way. If we additionally make the very weak assumption that at least one of the distinct policies in each $k$ iteration has a distinct base return from the others (as is generally likely to be the case), the inequality in the theorem above will be strict, i.e. networked learning will always be faster in expectation.
Proof.
Let us consider two scenarios. Firstly let us imagine that within the communication round, agents swap policies, but no policy drops out of the population, such that if agent $i$ adopts policy $\pi^{j}_{k+1}$, there exists an agent $i^{\prime}$ that adopts policy $\pi^{i}_{k+1}$, and so on. That way we end up with the same policies in the population as before the change, but with each one possibly carried by different arbitrary agents. This is equivalent to if no communication had taken place, meaning that in this scenario $V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})=V^{pop}(\bm{\pi}^{\mathrm{ind}}_{k+1},\mu_{t})$.
Let us now consider an alternative scenario. The softmax adoption scheme (Line 11, Alg. 3), which for $\tau^{comm}_{k}\in\mathbb{R}_{>0}$ gives non-uniform adoption probabilities for distinct $\sigma$ values, means by definition that some communicated policies are more likely to be adopted than others if they have distinct finitely estimated returns. Thus in expectation the number of distinct policies in the population will decrease if at least one policy has a distinct return from the others (of course, there is a possibility of this still happening even if no policy has a distinct return from the others). Let us start by saying for simplicity that during the first communication round a single $\pi^{(j,\mathrm{net})}_{k+1,c}$ is replaced by $\pi^{(i,\mathrm{net})}_{k+1,c}$, such that for $c=0$
$\bm{\pi}^{\mathrm{net}}_{k+1,c}$ $=\left(\pi^{(1,\mathrm{net})}_{k+1,c},\dots,\pi^{(\textbf{i},\mathrm{net})}_{k+1,c},\dots,\pi^{(\textbf{j},\mathrm{net})}_{k+1,c},\dots\pi^{(N,\mathrm{net})}_{k+1,c}\right)$,
$\text{and}\quad\bm{\pi}^{\mathrm{net}}_{k+1,c+1}$ $=\left(\pi^{(1,\mathrm{net})}_{k+1,c+1},\dots,\pi^{(\textbf{i},\mathrm{net})}_{k+1,c+1},\dots,\pi^{(\textbf{i},\mathrm{net})}_{k+1,c+1},\dots\pi^{(N,\mathrm{net})}_{k+1,c+1}\right)$.
For this to have occurred, we know that
$\mathbb{E}[\sigma^{(i,\mathrm{net})}_{k+1,c}]>\mathbb{E}[\sigma^{(j,\mathrm{net})}_{k+1,c}$],
and therefore by Assumption 5.6 that
$\mathbb{E}\left[V^{(i,\mathrm{net})}(\bm{\pi}^{\mathrm{net}}_{k+1,c},\mu_{t})\right]>\mathbb{E}\left[V^{(j,\mathrm{net})}(\bm{\pi}^{\mathrm{net}}_{k+1,c},\mu_{t})\right].$ (6)
By Assumption 5.5 we know that straight after the random policy updates there is no alignment among policies, i.e. in a coordination game we have $f_{c}^{(i,\mathrm{net})}$ = $f_{c}^{(j,\mathrm{net})}=\min f_{c}$. Therefore if Eq. 6 pertains, by Def. 5.2 it must be because:
i.e. because the base policy quality is higher for $\pi^{(i,\mathrm{net})}$ than for $\pi^{(j,\mathrm{net})}$. For this reason we have, for $c=0$:
$\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})\right]>\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c},\mu_{t})\right].$ (8)
Additionally, replacing $\pi^{(j,\mathrm{net})}_{k+1,c}$ with a second copy of $\pi^{(i,\mathrm{net})}_{k+1,c}$ will increase the alignment ($f_{c}$) of $\pi^{(i,\mathrm{net})}_{k+1,c}$ such that $\mathbb{E}\left[V^{(i,\mathrm{net})}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})\right]>\mathbb{E}\left[V^{(i,\mathrm{net})}(\bm{\pi}^{\mathrm{net}}_{k+1,c},\mu_{t})\right]$, increasing the improvement even further. This effect is even greater if more than one policy is replaced.
Since the independent case is equivalent to the networked case when $C_{p}=0$, we can say that $\bm{\pi}^{\mathrm{ind}}_{k+1}=\bm{\pi}^{\mathrm{net}}_{k+1,0}$. This gives the result, i.e.
$\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})\right]\geq\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{ind}}_{k+1},\mu_{t})\right],$
where this inequality is only not strict if the first scenario always applies rather than the second. ∎
To prove the benefit of the networked case over the independent case in anti-coordination games, we use a final additional assumption. This presumes that the base return is not yet fully maximised, and that the benefit to an agent’s overall return of increasing its base return by adopting a neighbour’s better-performing policy outweighs the resulting decrease in diversity. This establishes the conditions under which our policy adoption scheme is able to advantage networked agents over those whose policies are always independent. This assumption applies in most non-trivial scenarios (at least at the beginning of training), namely where the goal of the task is not simply for agents to have distinct policies that are otherwise inconsequential, and thus where the benefit of diverse behaviour can only be fully felt once agents have a certain level of aptitude at accomplishing the given task. For example, in all of the anti-coordination games in our experiments, agents are always penalised for moving, and only start to receive higher rewards if they are stationary. Therefore in these anti-coordination games agents will receive higher returns by aligning on policies that prioritise stationarity, than by maintaining diverse policies that have high levels of movement. Of course once base return is maximised and the assumption no longer holds, one can consider terminating policy communication and adoption to avoid decreases in diversity (one may also be ready to stop training entirely at this point, as the population is likely to be reaching the optimal average return). Please see Sec. 6.3 for further discussion of the applicability of this assumption in practice.
Assumption 5.14.
@@
Assume that an increase in the base return function outweighs a decrease in the policy diversity, namely $h(b+\Delta b,f_{d}-\Delta f_{d})>h(b,f_{d}),\;\forall\Delta b>0,\Delta f_{d}>0$, and that the agents have not yet maximised their base return function i.e. $b(\pi^{i}_{k+1})<\sup_{\pi\in\Pi}b(\pi)\;\;\forall i\in\{1,\dots,N\}$.
We now give our final theorem, namely that in anti-coordination games, in expectation networked populations will increase their returns at least as fast as independent ones with only a single round of communication in each iteration.
Theorem 5.15.
Let us once again set $\tau^{comm}_{k}\in\mathbb{R}_{>0}$. In an anti-coordination game, given Assumptions 5.5, 5.6, 5.11and 5.14, for $c=0$, $\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})\right]\geq\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{ind}}_{k+1},\mu_{t})\right]$.
Proof.
The proof begins similarly to that for a coordination game. Let us consider two scenarios. Firstly let us imagine that within the communication round, agents swap policies, but no policy drops out of the population, such that if agent $i$ adopts policy $\pi^{j}_{k+1}$, there exists an agent $i^{\prime}$ that adopts policy $\pi^{i}_{k+1}$, and so on. That way we end up with the same policies in the population as before the change, but with each one possibly carried by different arbitrary agents. This is equivalent to if no communication had taken place, meaning that in this scenario $V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})=V^{pop}(\bm{\pi}^{\mathrm{ind}}_{k+1},\mu_{t})$.
Let us now consider an alternative scenario. The softmax adoption scheme (Line 11, Alg. 3), which for $\tau^{comm}_{k}\in\mathbb{R}_{>0}$ gives non-uniform adoption probabilities for distinct $\sigma$ values, means by definition that some communicated policies are more likely to be adopted than others if they have distinct finitely estimated returns. Thus in expectation the number of distinct policies in the population will decrease if at least one policy has a distinct return from the others. Say for simplicity that during the first communication round a $\pi^{(j,\mathrm{net})}_{k+1,c}$ is replaced by $\pi^{(i,\mathrm{net})}_{k+1,c}$, such that for $c=0$
$\bm{\pi}^{\mathrm{net}}_{k+1,c}$ $=\left(\pi^{(1,\mathrm{net})}_{k+1,c},\dots,\pi^{(\textbf{i},\mathrm{net})}_{k+1,c},\dots,\pi^{(\textbf{j},\mathrm{net})}_{k+1,c},\dots\pi^{(N,\mathrm{net})}_{k+1,c}\right)$,
$\text{and}\quad\bm{\pi}^{\mathrm{net}}_{k+1,c+1}$ $=\left(\pi^{(1,\mathrm{net})}_{k+1,c+1},\dots,\pi^{(\textbf{i},\mathrm{net})}_{k+1,c+1},\dots,\pi^{(\textbf{i},\mathrm{net})}_{k+1,c+1},\dots\pi^{(N,\mathrm{net})}_{k+1,c+1}\right)$.
For this to have occurred, we know that
$\mathbb{E}[\sigma^{(i,\mathrm{net})}_{k+1,c}]>\mathbb{E}[\sigma^{(j,\mathrm{net})}_{k+1,c}$],
and therefore by Assumption 5.6 that
$\mathbb{E}\left[V^{(i,\mathrm{net})}(\bm{\pi}^{\mathrm{net}}_{k+1,c},\mu_{t})\right]>\mathbb{E}\left[V^{(j,\mathrm{net})}(\bm{\pi}^{\mathrm{net}}_{k+1,c},\mu_{t})\right].$ (9)
By Assumption 5.5 we know that straight after the random policy updates there is no alignment among policies, i.e. in the anti-coordination game we have $f_{d}^{(i,\mathrm{net})}$ = $f_{d}^{(j,\mathrm{net})}=\max f_{d}$, while by Assumption 5.14 we know that the agents have not yet maximised their base return function. Therefore if Eq. 9 pertains, by Def. 5.2 it must be because:
Assumption 5.14 assumes that any increase in the base quality of the policy will outweigh the decrease in diversity that will come from having more than one agent following $\pi^{(i,\mathrm{net})}_{k+1,c+1}$. Therefore we have, for $c=0$:
$\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})\right]>\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c},\mu_{t})\right].$
These steps apply similarly if more than one policy is replaced.
Since the independent case is equivalent to the networked case when $C_{p}=0$, we can say that $\bm{\pi}^{\mathrm{ind}}_{k+1}=\bm{\pi}^{\mathrm{net}}_{k+1,0}$. This gives the result, i.e.
$\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{net}}_{k+1,c+1},\mu_{t})\right]\geq\mathbb{E}\left[V^{pop}(\bm{\pi}^{\mathrm{ind}}_{k+1},\mu_{t})\right],$
where this inequality is only not strict if the first scenario always applies rather than the second. ∎
@@ 6.1 Experimental setup
We present experiments from grid worlds, following the gold standard in similar works on MFGs and MFC (survey_learningMFGs). We give results from six tasks similar to those found in prior works, defined by the agents’ reward/transition functions and relating to agents’ positions relative to other agents. Two are coordination games and four are anti-coordination games, where in each case the reward function reflects a coordination/anti-coordination ($f_{c}/f_{d}$) element alongside other elements that may be crucial for receiving reward, reflected in the policies’ base quality $b(\pi)$ (Sec. 5). In all cases, rewards are normalised in [0,1] after they are computed.
The two coordination games are:
[…]
  • Target selection. This game is also used in benjamin2024networked; benjamin2024networkedapproximation. Unlike in the above ‘cluster’ game, the agents are given options of locations at which to gather, and they must reach consensus among themselves. If the agents are co-located with one of a number of specified targets $\phi\in\Phi$ (in our experiments we place one target in each of the four corners of the grid), and other agents are also at that target, they get a reward proportional to the fraction of the population found there; otherwise they receive a penalty of -1. In other words, the agents must coordinate on which of a number of mutually beneficial points will be their single gathering place. Define the magnitude of the distances between $x,y$ at $t$ as $dist_{t}(x,y)$. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{targ}(r_{coord}(\hat{\mu}_{t}(s^{i}_{t})))$, where
[…]
@@
  • Disperse. This game is also used in benjamin2024networkedapproximation and is similar to the ‘exploration’ tasks in scalable_deep; wu2024populationaware and other MFG works. In our version agents are rewarded for being located in more sparsely populated areas but only if they are stationary, to avoid trivial random policies. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{stationary}(-\log(\hat{\mu}_{t}(s^{i}_{t})))$, where
[…]
  • Target coverage. The population is rewarded for spreading across a certain number of targets, as long as agents are stationary at the target. As in the ‘target selection’ game, we have targets $\phi\in\Phi$, where in our experiments we place one target in each of the four corners of the grid. Again define the magnitude of the distances between $x,y$ at $t$ as $dist_{t}(x,y)$. The reward function is given by
    where $r_{stationary}$ and $r_{targ}$ are as defined above.
  • Beach bar. Such games are very common in MFG works (continuous_fictitious_play; survey_learningMFGs; cui2023multiagent; wu2024populationaware). In our version agents are rewarded for being stationary in sparsely populated locations as close as possible to a target $\phi_{b}$, located in the centre of the grid. The maximum possible distance from the target is denoted maxDist. The reward is given by
    where $r_{stationary}$ is as defined above.
  • Shape formation. The population is rewarded for spreading around a ring shape, accomplished by encouraging agents to be a distance of 3 (chosen arbitrarily to fit the grid) from a centre point $\phi_{c}$. The reward is given by
    where $r_{stationary}$ is as defined above, and
[…]
In these spatial environments, we define both the communication network $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$ by the physical distance from $i$. We show plots for various transmission radii, given as fractions of the maximum distance in the grid. Note that the networked population with the largest radius is always fully connected, and therefore these agents are always able to accurately estimate $\hat{r}$ and ${\hat{\mu}}_{t}$ even for $C_{r}=1$ and $C_{e}=0$. That is, when we set $C_{r}=C_{e}>0$ their observations are equivalent to those that the central-agent population would receive, albeit that policies are updated and spread differently.
We evaluate our experiments according to a finite-step estimation of the population-average discounted return (Def. 3.3) over $M$ steps within each outer $k$ loop, i.e. $\hat{V}^{pop}(\bm{\pi}_{k},\mu_{t};M)$. Experiments were conducted on a Linux-based machine with 2 x Intel Xeon Gold 6248 CPUs (40 physical cores, 80 threads total, 55 MiB L3 cache). We use the JAX framework to accelerate and vectorise our code. We run five trials with different random seeds for each experiment, and plot the mean and standard deviation of the mean across the seeds. Random seeds are set in our code in a fixed way dependent on the trial number to allow easy replication of experiments. Our code is included in the publicly available supplementary material for reproducibility.
6.2 Hyperparameters
Table 1: Hyperparameters
Hyperparam. Value Comment
Trials 5 We run 5 trials with different random seeds for each experiment. We plot the mean and standard deviation of the mean for each metric across the seeds.
Gridsize 20x20 -
Population 500 We chose 500 for our demonstrations to show that our algorithm can handle large populations, indeed often larger than those demonstrated in other mean-field works, especially for grid-world environments, while also being feasible to simulate with respect to time and computation constraints (mfrl_yang; rl_stationary; subramanian2022multi; pomfrl; approximately_entropy; subjective_equilibria; DMFG; cui2023multiagent; general_framework; benjamin2024networked; benjamin2024networkedapproximation; wu2024populationaware). For example, the MFC work in carmona2019linear uses 10 agents; the work on decentralised execution for MFC by cui2023learningdecentralisedcontrol uses 200 agents.
Number of neurons in input layer 440 The agent’s position is represented by two concatenated one-hot vectors, indicating the agent’s row and column. The mean-field distribution is a flattened vector of the same size as the grid. As such, the input size is $(2\times\mathrm{dimension})+(\mathrm{dimension}^{2})]$.
Hidden layers 2 We achieved sufficient learning speed with just 2 hidden layers, but further optimising the number of layers may lead to better results.
Activation function ReLU This is a common choice in deep RL.
$C_{p}$ 1 (10/50) As in benjamin2024networked; benjamin2024networkedapproximation, we choose a value of 1 for most experiments to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity. We also conduct additional studies to show the effect of additional rounds in Figs. 4 and 5.
$C_{r}$ 1 (10/50) Similar to $C_{p}$, we choose this value to show our algorithm’s ability to appropriately estimate the average reward even with only a single round, even in networks that may have limited connectivity. We conduct additional studies to show the effect of additional rounds in Figs. 4 and 5.
$C_{e}$ 1 (10/50) Similar to $C_{p}$, we choose this value to show the ability of our algorithm to appropriately estimate the mean field even with only a single communication round, even in networks that may have limited connectivity. We also conduct additional studies to show the effect of additional rounds in Figs. 4 and 5.
$\gamma$ 0.9 Standard choice across RL literature.
$\tau_{q}$ 0.03 We follow NEURIPS2020_2c6a0bae and benjamin2024networkedapproximation, which tested a range of values.
clcl -1 We use the same value as in NEURIPS2020_2c6a0bae and benjamin2024networkedapproximation.
$\nu$ L1L-1 We follow benjamin2024networkedapproximation, which is similar to scalable_deep.
Optimiser Adam As in NEURIPS2020_2c6a0bae, we use the Adam optimiser with initial learning rate 0.01.
$\tau^{comm}_{k}$ cf. comment We follow benjamin2024networkedapproximation, where $\tau^{comm}_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations. Further optimising this inverse annealing process may lead to better results; we provide an ablation study in Fig. 10.
In our experiments we generally want to demonstrate that our communication-based algorithm learns faster than the central-agent and independent architectures, even when the Q-function / mean field / average reward are poorly estimated as is likely to be the case in complex real-world scenarios. There is a similar motivation in the related works on networked communication in the MFG setting by benjamin2024networked; benjamin2024networkedapproximation. Moreover we want to show that there is a benefit even to a small amount of communication, so that communication rounds themselves do not excessively add to time complexity. As such, we generally select hyperparameters at the lowest end of those we tested during development, to show that our algorithms are particularly successful and robust given what might otherwise be considered ‘undesirable’ hyperparameter choices.
6.3 Results and discussion